### 17 Mar 2026 v2

# Grassroots Bonds: A Grassroots Foundation for Market Liquidity

Abstract.
Global cryptocurrencies are unbacked and have high transaction cost incurred by global consensus. In contrast, grassroots cryptocurrencies are backed by the goods and services of their issuers—any person, natural or legal—and have no transaction cost beyond operating a smartphone. Liquidity in grassroots cryptocurrencies arises from mutual credit via coin exchange among issuers. However, as grassroots coins are redeemable 1-for-1 against any other grassroots coin, the credit-forming exchange must also be 1-for-1, lest prompt redemption after exchange would leave the parties with undue profit or loss. Thus, grassroots coins are incongruent with liquidity through interest-bearing credit.
Here we introduce grassroots bonds, which extend grassroots coins with a maturity date, reframing grassroots coins—cash—as mature grassroots bonds. Bond redemption generalises coin redemption, allowing the lending of liquid coins in exchange for interest-bearing future-maturity bonds. We show that digital social contracts—voluntary agreements among persons, specified, fulfilled, and enforced digitally—can express the full gamut of financial instruments as the voluntary swap of grassroots bonds, including credit lines, loans, sale of debt, forward contracts, options, and escrow-based instruments, and that classical liquidity ratios are applicable just as well to grassroots bonds. Grassroots bonds may thus allow local digital economies to form and grow without initial capital or external credit, harnessing mutual trust within communities into liquidity.
The formal specification presented here was used by AI to derive a working implementation of grassroots bonds in GLP, a concurrent logic programming language implemented in Dart for smartphone deployment. The implementation is illustrated by a running multiagent village market scenario, also implemented in GLP by AI.
1. Introduction
Grassroots currencies. Grassroots coins (Shapiro, 2024b) are units of debt that can be issued and traded digitally by any person—natural or legal—including people, communities, cooperatives, corporations, banks, municipalities and governments. Grassroots currencies are more similar to fiat currencies in that they are backed by goods and services offered by the issuer, who controls their scarcity, as well as to ‘inside money’ (Cavalcanti and Wallace, 1999b, a)—a medium of exchange backed by private credit—than to global cryptocurrencies such as Bitcoin (Nakamoto and Bitcoin, 2008) or Ethereum (Buterin, 2014), which are unbacked and for which scarcity is controlled by the protocol.
Grassroots currencies aimed to provide a foundation for grassroots digital economies, allowing them to emerge without initial capital or external credit, operating solely on the networked smartphones of their members, and gradually merge into a global digital economy. As such, grassroots currencies may empower economically-deprived communities worldwide, and by harnessing mutual trust within communities help ‘banking the unbanked’ (Bruhn and Love, 2009; Agarwal et al., 2017). These abilities may be most relevant in places where capital and external credit are scarce but trust within families and communities is abundant.
Value and redemption. Grassroots currencies are endowed with value by their issuer undertaking to redeem them against their offerings: Goods and services for people and corporations; interest and transaction fees for banks; taxes for cities and states; and fiat currencies for central banks. Critically, the offerings of each person must also include all the grassroots coins it holds, including coins issued by other persons. Upon request, a person must redeem any coin it has issued, 1-for-1, against any coin it holds.
Coin redemption is key for a digital economy based on grassroots currencies to function, as it (i) makes coins issued by the same person fungible; (ii) resolves doublespending (Lewis-Pye et al., 2023); (iii) allows the revocation of credit; (iv) allows chain payments across mutually-liquid currencies; and (v) pegs mutually-liquid grassroots currencies at a 1-for-1 exchange rate.
Liquidity through mutual credit. As each person prices their goods and services in terms of their own currency, trade is impossible if no person holds another person’s coins. This can be remedied by mutual credit lines, formed by the voluntary exchange of grassroots coins among persons that know and trust each other—family members, friends, peers, and colleagues; a community bank and its members; a corporation and its owners, employees, suppliers, and customers. Liquidity arises from persons holding each other’s coins, and is correlated with the amount and distribution of coins in circulation. A grassroots coin in circulation is an outstanding obligation by its issuer to its holder, and the amount of grassroots coins in circulation in a digital economy is a measure of the trust among its members.
The limitation of coins. Since grassroots coins can be redeemed 1-for-1, mutual credit must also be 1-for-1, lest closing it right after formation would leave the parties with undue profit or loss. Concretely, if $p$ opens a mutual line of credit with $q$ with an upside for $p$, e.g. 100 $p$-coins against 110 $q$-coins, then coin redemption would allow $p$ to immediately redeem the 100 $q$-coins it holds, leaving $p$ with 10 $q$-coins and $q$ with nothing; this is akin to a bank granting a loan with an upfront interest payment then immediately recalling the loan while keeping the interest. More generally, grassroots currencies cannot associate interest with credit in an intrinsic way.
Grassroots bonds. Here we present grassroots bonds, which extend grassroots coins with a maturity date, reframing grassroots coins—cash—as mature grassroots bonds. Coin-for-bond redemption generalises coin-for-coin redemption, allowing the lending of liquid coins in exchange for interest-bearing future-maturity bonds, facilitating incentive-based liquidity.
Implementation. The grassroots bonds specification presented here was used by AI to derive a working implementation in GLP (Shapiro, 2025), a multiagent concurrent logic programming language implemented in Dart for smartphone deployment. The implementation is demonstrated by a running six-agent village market scenario, also implemented in GLP by AI, that exercises symmetric mutual credit, asymmetric credit with interest, payments, portfolio swaps, escrow, redemption, and sale of debt.
Security. A secure grassroots payment system for grassroots coins has been presented in (Lewis-Pye et al., 2023). Its security is based on the issuer of each currency—its sovereign—guaranteeing the security of transactions in that currency: the sovereign approves or disapproves payments in their coins, and safety violations are detectable as signed cryptographic evidence. No consensus is needed; a sovereign that misbehaves or fails to respond risks their coins being treated as bad debt. This model applies directly to unilateral bond transactions (mint, pay, redeem); voluntary swaps between two parties employ bilateral atomic commitment, with escrow (Section 4) as a non-blocking alternative. Integrating these mechanisms in a secure implementation of GLP is a subject of ongoing work.
Contributions and paper structure. Section 2 presents grassroots bonds and their specification as a digital social contract with two transactions—mint and swap—from which all financial instruments emerge. Section 3 develops the full gamut of financial instruments as volitional bond swaps, including credit lines, loans, sale of debt, forward contracts, and options. Section 4 adds escrow-based instruments: collateral, guarantees, options, insurance, credit default swaps, and letters of credit. Section 5 applies classical liquidity ratios—cash, quick, and current—to grassroots bonds. Section 6 develops a community bank scenario tying the instruments together. Section 7 implements the bonds social contract in GLP (Shapiro, 2025), a multiagent concurrent logic programming language, demonstrated in a six-agent village market scenario. Section 8 compares with global cryptocurrencies, DeFi and classical finance. Section 9 concludes. Appendix A provides mathematical foundations for the work presented here. Appendix B presents the bond agent GLP source code.
2. Grassroots Bonds
We assume a global present time $t^{*}\in\mathbb{N}$ and a potentially infinite set of agents $\Pi$, but consider only finite subsets of it, so when we refer to a particular set of agents $P\subset\Pi$ we assume $P$ to be nonempty and finite.
Definition 2.1 (Grassroots Bonds).
A $p$-bond with maturity date $t$, denoted ¢p,t, is a unit of debt issued by $p\in\Pi$ maturing at time $t\in\mathbb{N}$; it is mature if $t\leq t^{*}$, in which case it is also referred to as a $p$-coin and denoted ¢p. In particular, a bond minted with $t=0$ is a grassroots coin (cf. (Shapiro, 2024b)) from the outset. We let $\mathcal{B}(P)=\{$$_{p,t}:p\in P,t\in\mathbb{N}\}$ denote the set of all grassroots bonds by agents $P\subset\Pi$, and write $^{k}_{p,t}$ for a multiset of $k$ bonds ¢p,t.
In a practical implementation, each $p$-bond will have a serial number and value and be signed by $p$. For simplicity of exposition the bonds (and coins) employed here can form multisets and are implicitly with denomination $1$.
As discussed in (Shapiro, 2024b), coins are backed by goods and services offered by the coin’s issuer and priced in their coins, with bond values derived from expectation on the coin value. Coin redemption ensures that coins issued by mutually-liquid persons are of equal value, namely traded at a 1-for-1 rate. A consequence of the redemption rule is that all coins held by an issuer of a coin are necessarily part of the goods offered by the issuer, priced 1-for-1. Thus, the holder of a $q$-coin that loses trust in $q$ may redeem it against any bond held by $q$. A holder of an immature $q$-bond who loses trust in $q$ may sell the debt in the free market, as described in Section 3.
2.1. A Digital Social Contract for Grassroots Bonds
A digital social contract (Cardelli et al., 2020) is a voluntary agreement among persons, specified, fulfilled, and enforced digitally. It is the grassroots counterpart of smart contracts (Shapiro, 2023), executed not by third-parties but by the parties to the contract, and may be consensus-based (Keidar et al., 2025) or consensus-free. Here, we specify the digital social contract that governs grassroots bonds using volitional multiagent atomic transactions, defined below, and demonstrate their implementation in Section 7.
Volitional transactions extend multiagent atomic transactions (Shapiro, 2026) by assuming that each agent participating in a transaction consists of a person operating a machine (e.g., smartphone), and encoding the persons that need to consent to a transaction. This requires a correct implementation of a volitional transaction to obtain such consent, for example using the person-machine interface. Note that personal consent is in addition to the basic requirement that all machines participating in an atomic transaction agree to it (Gray, 1981), even in transactions that the consent of the person operating them is not needed.
We use $S^{P}$ to denote the set $S$ indexed by $P$, with $c_{p}$ the component indexed by $p$.
Definition 2.2 (Configuration, Transaction, Volitional Transaction).
A configuration over agents $P\subset\Pi$ and a set of local states $S$ is a member $c\in S^{P}$. A transaction over $P$ and $S$ is a pair $t=c\rightarrow c^{\prime}$ of distinct configurations, each over $P$ and $S$; agent $p$ is an active participant in $t$ if $c_{p}\neq c^{\prime}_{p}$, stationary otherwise. A volitional transaction is a pair $(t,V)$ where $t$ is such a transaction and $V\subseteq P$ is the set of volitional participants—those whose person must consent for the transaction to execute. In such a case we say that the transition $t$ is guarded by $V$.
Appendix A provides mathematical background, in particular how a set of multiagent atomic transactions induces a family of multiagent transition systems (Shapiro, 2026) and the conditions under which they are grassroots (Shapiro, 2023).
The grassroots bonds social contract is specified by two volitional transactions: Mint, guarded by its sole participating agent, by which the agent mints new bonds; and Swap, by which its two participating agents $p$ and $q$ swap bonds. The latter has three sub-cases: Voluntary swap, by which the two agents agree to swap two sets of bonds, guarded by both $p$ and $q$; Pay, by which $p$ unilaterally pays $q$ with $q$-coins, guarded by $p$; and Redeem, by which $p$ unilaterally transfers to $q$ a $q$-coin in exchange for a bond held by $q$, guarded by $p$.
Only coins—that is, mature bonds—may be paid or redeemed; coin-for-bond redemption generalises the coin-for-coin redemption of grassroots currencies (Shapiro, 2024b) by allowing the holder of a coin to choose any bond held by the issuer of the coin. The holder of an immature bond may sell the debt to a third party, as shown in Section 3.
Definition 2.3 (Grassroots Bonds Volitional Transactions).
For agents $P\subset\Pi$, the set of states $S(P)$ consists of all multisets of members of $\mathcal{B}(P)$, with initial state $s_{0}=\emptyset$. The grassroots bonds volitional transactions over $P$ and $S(P)$ are all transactions such that:
  1. (1)
    Mint: A unary $p$-transaction $s\rightarrow s^{\prime}$, $p\in P$, guarded by $p$, $s^{\prime}:=s\cup\{\text{\textcent}^{k}_{p,t}\}$, $k,t\in\mathbb{N}$.
  2. (2)
    Swap: A binary $\{p,q\}$-transaction, $c\rightarrow c^{\prime}$, $\{p,q\}\subseteq P$, $c^{\prime}_{p}:=(c_{p}\cup y)\setminus x$, $c^{\prime}_{q}:=(c_{q}\cup x)\setminus y$ for any $x,y$ for which one of the following holds:
    1. (a)
      Voluntary swap: $x\subseteq c_{p}$, $y\subseteq c_{q}$. Guarded by $\{p,q\}$.
    2. (b)
      Pay: $x\subseteq c_{p}$ is a set of $q$-coins, $y=\emptyset$. Guarded by $p$.
    3. (c)
      Redeem: $x=\{\text{\textcent}_{q}\}\subseteq c_{p}$, $y=\{\text{\textcent}_{r,t}\}\subseteq c_{q}$, $t\in\mathbb{N}$. Guarded by $p$.
We note four properties of grassroots bonds:
  1. (1)
    Conservation of Money: In a run of a grassroots bonds transition system, the $p$-bonds in any configuration $c$ in the run are exactly the $p$-bonds minted by $p$ in the prefix of the run ending in $c$.
  2. (2)
    Chain redemption: Coin redemption generalises the chain redemptions of grassroots coins (Shapiro, 2024b). If there is 1-liquidity from $p$ to $q$, defined by the existence of a sequence of persons $p_{0},\ldots,p_{k}$, $p_{0}=p$, $p_{k}=q$, where each $p_{i}$, $0\leq i<k$, holds a $p_{i+1}$-coin, then $p$ can initiate a chain of k1k-1 redemptions to obtain a $q$-coin. Chain redemption enables long-range payments and arbitrage, and entails a 1:1 exchange rate among coins within liquidity-connected components.
  3. (3)
    Insolvency: Insolvency is manifest in the inability of a person to satisfy redemption claims on coins they have issued. Holders of an insolvent person’s bonds may treat them as bad debt and sell them at a discount, effectively devaluating the insolvent person’s currency (Shapiro, 2024b). The instruments developed in this paper—collateral, escrow, and community-level risk management (Sections 46)—provide measures to mitigate the consequences of insolvency.
  4. (4)
    Grassroots: The grassroots bonds protocol is grassroots (Shapiro, 2023) in the formal sense of (Shapiro, 2026): multiple independent deployments can emerge without global resources and interoperate once interconnected. The proof is given in Appendix A.
3. Financial Instruments
The full gamut of financial instruments can be expressed via the voluntary swap of grassroots bonds. Each instrument below is a voluntary swap (Definition 2.3) between agents $p$ and $q$, guarded by the two agents, specified by the bonds $x$ transferred from $p$ to $q$ and the bonds $y$ transferred from $q$ to $p$.
  1. (1)
    Symmetric mutual credit. $p$ and $q$ each mint $k$ coins and exchange them: $x=\text{\textcent}^{k}_{p}$, $y=\text{\textcent}^{k}_{q}$.
  2. (2)
    Zero-coupon loan. Lender $p$ mints $k^{\prime}<k$ coins, borrower $q$ mints a bond maturing at $t$: $x=\text{\textcent}^{k^{\prime}}_{p}$, $y=\text{\textcent}^{k}_{q,t}$.
  3. (3)
    Balloon loan. Lender $p$ mints $k$ coins, borrower $q$ mints interest bonds $\text{\textcent}^{k_{j}}_{q,t_{j}}$ for $j\in[1..n]$ and a principal bond $\text{\textcent}^{k}_{q,t}$: $x=\text{\textcent}^{k}_{p}$, $y=\bigcup_{j=1}^{n}\text{\textcent}^{k_{j}}_{q,t_{j}}\cup\text{\textcent}^{k}_{q,t}$.
  4. (4)
    Fixed-payment loan. Lender $p$ mints $k$ coins, borrower $q$ mints bonds with fixed payments $k_{j}$ at dates $t_{j}$: $x=\text{\textcent}^{k}_{p}$, $y=\bigcup_{j=1}^{n}\text{\textcent}^{k_{j}}_{q,t_{j}}$.
  5. (5)
    Revolving credit line. An asymmetric credit line from $p$ to $q$ with limit $k$: $x=\text{\textcent}^{k}_{p}$, $y=\text{\textcent}^{k}_{q,t}$. When $q$ repays (pays $p$ in $p$-coins), $q$ may draw again: a new voluntary swap up to the remaining limit, consented to by $p$.
  6. (6)
    Sale of debt. $p$ sells $r$-bonds to $q$ for $k^{\prime}<k$ $q$-coins: $x=\text{\textcent}^{k}_{r}$, $y=\text{\textcent}^{k^{\prime}}_{q}$. The debtor $r$ is not a party.
  7. (7)
    Forward contract. $p$ and $q$ exchange bonds both maturing at $t$: $x=\text{\textcent}^{k}_{p,t}$, $y=\text{\textcent}^{k^{\prime}}_{q,t}$.
  8. (8)
    Interest rate swap. $p$ and $q$ exchange two sets of bonds with differing maturity schedules: $x=\bigcup_{j}\text{\textcent}^{k_{j}}_{p,t_{j}}$, $y=\bigcup_{j}\text{\textcent}^{k^{\prime}_{j}}_{q,t_{j}}$, where $p$ pays fixed amounts $k_{j}$ while $q$ pays amounts $k^{\prime}_{j}$ that vary according to a reference rate. Periodic settlement swaps realise the net difference at each date.
The instruments cover mutual credit and loans (items 1–5), debt trading (item 6), derivatives (items 7–8). Currency swaps and repurchase agreements (repos) are realised as sequences of two swaps; factoring (selling receivables at a discount) is an instance of sale of debt; a mortgage is a loan (items 2–4) combined with collateral. Collateral, guarantees, options, insurance, and further escrow-based instruments are developed next.
4. Escrow
An escrow agent $e$ holds bonds on behalf of others and releases them according to agreed-upon conditions. No new transaction type is needed: deposit, release, and return are all swaps, guarded by the parties to the swap.
Mechanism. An escrow arrangement involves a depositor, a beneficiary, and an escrow agent $e$. The depositor transfers bonds to $e$ via a swap. Subsequently, $e$ either releases the bonds to the beneficiary or returns them to the depositor, each via a swap. The conditions under which $e$ releases or returns are part of the agreement.
  1. (1)
    Collateral: To collateralise a loan from lender $p$ to borrower $q$:
    1. (a)
      Deposit. $\{(q,x),\;(e,\emptyset)\}$: borrower $q$ transfers bonds $x$ to escrow agent $e$.
    2. (b)
      Release on default. $\{(e,x),\;(p,\emptyset)\}$: $e$ transfers collateral to lender $p$.
    3. (c)
      Return on fulfilment. $\{(e,x),\;(q,\emptyset)\}$: $e$ returns collateral to borrower $q$.
  2. (2)
    Guarantee: Guarantor $g$ covers borrower $q$’s obligations to lender $p$:
    1. (a)
      Deposit. $\{(g,\text{\textcent}^{k}_{g}),\;(e,\emptyset)\}$: $g$ deposits $g$-bonds with escrow agent $e$.
    2. (b)
      Invocation on default. $\{(e,\text{\textcent}^{k}_{g}),\;(p,\emptyset)\}$: $e$ releases $g$-bonds to lender $p$.
    3. (c)
      Release on fulfilment. $\{(e,\text{\textcent}^{k}_{g}),\;(g,\emptyset)\}$: $e$ returns $g$-bonds to $g$.
  3. (3)
    Option: An option grants $p$ the right, but not the obligation, to execute a specified swap with $q$ at or before date $t$. The option premium and the underlying assets are held in escrow:
    1. (a)
      Establishment. $p$ deposits the premium $^{k}_{p}$ with escrow agent $e$: $\{(p,\text{\textcent}^{k}_{p}),\;(e,\emptyset)\}$. $q$ deposits the underlying bonds $y$ with $e$: $\{(q,y),\;(e,\emptyset)\}$.
    2. (b)
      Exercise. $p$ signals exercise. $e$ releases the underlying $y$ to $p$: $\{(e,y),\;(p,\emptyset)\}$, and releases the premium to $q$: $\{(e,\text{\textcent}^{k}_{p}),\;(q,\emptyset)\}$.
    3. (c)
      Expiry. If $p$ has not exercised by $t$, $e$ returns the underlying to $q$: $\{(e,y),\;(q,\emptyset)\}$, and transfers the premium to $q$: $\{(e,\text{\textcent}^{k}_{p}),\;(q,\emptyset)\}$.
    The above specifies a call option: $p$ acquires the right to buy the underlying $y$ from $q$. A put option—the right to sell bonds to $q$ at a predetermined price—is symmetric: $q$ deposits the strike price with $e$, and $p$ deposits the bonds it wishes to sell.
    The specification above uses American exercise: $p$ may exercise at or before $t$. For European exercise, $p$ may exercise only at $t$.
    A strike price can be incorporated by having $p$ deposit additional bonds $^{k^{\prime}}_{p}$ with $e$ at establishment. On exercise, $e$ transfers the strike price to $q$ along with the premium; on expiry, $e$ returns the strike price to $p$ along with the premium.
  4. (4)
    Insurance: Party $p$ (the insured) pays a premium to escrow agent $e$, who disburses a payout if a specified event occurs, as attested by $e$:
    1. (a)
      Premium. $\{(p,\text{\textcent}^{k}_{p}),\;(e,\emptyset)\}$: $p$ deposits the premium with $e$.
    2. (b)
      Claim (event occurs). $\{(e,\text{\textcent}^{k^{\prime}}_{q}),\;(p,\emptyset)\}$: $e$ transfers the payout $k^{\prime}\gg k$ to $p$. The payout is funded by the insurer $q$, who has deposited bonds with $e$ as reserves.
    3. (c)
      Expiry (no event). The premium is retained by $e$ on behalf of insurer $q$.
    The escrow agent $e$ serves as the claims adjudicator. A mutual insurance arrangement among a group of participants can be realised by each member depositing premiums into a shared escrow reserve.
  5. (5)
    Credit default swap (CDS): Party $p$ pays periodic premiums to $q$ (the protection seller), and $q$ compensates $p$ if a reference entity $r$ defaults. Escrow agent $e$ adjudicates:
    1. (a)
      Premium payments. Periodic swaps $\{(p,\text{\textcent}^{k_{j}}_{p,t_{j}}),\;(e,\emptyset)\}$: $p$ deposits premium bonds with $e$, forwarded to $q$.
    2. (b)
      Credit event. $\{(e,\text{\textcent}^{K}_{q}),\;(p,\emptyset)\}$: if $e$ determines that $r$ has defaulted, $e$ releases $q$’s reserve bonds to $p$.
    3. (c)
      No credit event. Premium bonds are released to $q$ and reserve bonds returned to $q$ at maturity.
  6. (6)
    Letter of credit: A bank $b$ guarantees payment to seller $p$ on behalf of buyer $q$, with escrow agent $e$ verifying that contractual conditions (e.g. delivery of goods) are met:
    1. (a)
      Issuance. $\{(b,\text{\textcent}^{k}_{b}),\;(e,\emptyset)\}$: bank $b$ deposits $b$-coins with $e$.
    2. (b)
      Presentation. $\{(e,\text{\textcent}^{k}_{b}),\;(p,\emptyset)\}$: upon $e$ verifying that $p$ has fulfilled the terms (e.g. delivery attestation), $e$ releases $b$-coins to seller $p$.
    3. (c)
      Reimbursement. $\{(q,\text{\textcent}^{k}_{q,t}),\;(b,\emptyset)\}$: buyer $q$ reimburses bank $b$ with $q$-bonds.
5. Liquidity Measures
The liquidity measures of cash, quick, and current ratios were developed in corporate finance to assess the liquidity of corporations. They were adapted to grassroots cryptocurrencies in (Shapiro, 2024b). Here we extend them to grassroots bonds, where maturity dates play a central role.
Recall that $t^{*}$ denotes the current time. Let $\nu_{p,t}(q)$ be the number of $p$-bonds with maturity date $\leq t$ held by $q$ at present, where $p\neq q$ (zero if $p=q$, as $p$-bonds held by $p$ carry no value). In particular, $\nu_{p,t^{*}}(q)$ is the number of mature $p$-bonds held by $q$, and $\nu_{p,\infty}(q)$ is the total number of $p$-bonds held by $q$.
Recall the redemption rule (Definition 2.3): a holder of a $p$-coin can redeem it against any bond held by $p$, chosen by the holder. Only coins—i.e., mature bonds—may be offered for redemption.
Cash ratio. The cash ratio measures the resilience of $p$ against a “run on the bank”: all holders of $p$-coins attempt to redeem at once. Under the redemption rule, each holder claims any bond from $p$’s holdings.
Let $\Delta(p,q):=\min(\nu_{p,t^{*}}(q),\;\nu_{q,t^{*}}(p))$. Thus $\Delta(p,q)$ is the number of mature $p$-bonds held by $q$ that $q$ can redeem with mature $q$-bonds held by $p$.
$\textit{Cash Ratio of }p=1-\frac{\sum_{q\in P}(\nu_{p,t^{*}}(q)-\Delta(p,q))}{\sum_{q\in P}\nu_{p,t^{*}}(q)}$
If $p$ holds enough mature foreign bonds to satisfy all redemption claims on its mature bonds, the cash ratio is 1. If $p$ holds no foreign bonds, the cash ratio is 0.
Quick ratio. Since only coins may be redeemed, the quick ratio considers all $p$-bonds in circulation as future liabilities that will become redeemable upon maturity, measured against all foreign bonds held by $p$ that will similarly mature into coins.
Let $\Delta_{\infty}(p,q):=\min(\nu_{p,\infty}(q),\;\nu_{q,\infty}(p))$.
$\textit{Quick Ratio of }p=1-\frac{\sum_{q\in P}(\nu_{p,\infty}(q)-\Delta_{\infty}(p,q))}{\sum_{q\in P}\nu_{p,\infty}(q)}$
The quick ratio may be smaller than the cash ratio if $p$ has many immature bonds in circulation relative to its holdings of foreign bonds.
Current ratio. The current ratio counts all foreign bonds held by $p$ as assets and all $p$-bonds in circulation as liabilities, without regard to redeemability or maturity matching:
$\textit{Current Ratio of }p=\frac{\sum_{q\in P}\nu_{q,\infty}(p)}{\sum_{q\in P}\nu_{p,\infty}(q)}$
The current ratio admits assets that may not be realisable: foreign bonds that cannot be redeemed, bonds by insolvent issuers, and fake assets issued by sybil persons to inflate the balance sheet of $p$.
6. Community Bank
A community bank, or credit union, may streamline liquidity and simplify payments within a community that employs grassroots bonds. The bank would normally have a governing body and signatories as authorised by the community. The community bank issues its own grassroots currency and may open credit lines with community members at its discretion (e.g., employing the objective measures for creditworthiness and liquidity discussed in Section 5), just as a private banker would do.
A key characteristic of a community bank is that its members undertake to accept payments in the community currency, in addition to their personal currency. Thus, transactions among liquid community members may include one-step payments in the community currency. Presumably, community members will be happy to use the community currency, knowing that they own the bank. Note that a community bank does not (necessarily) need initial capital, and may employ standard measures to increase its capital such as charging interest on drawn credit and transaction fees.
Community mutual credit scenario. Consider a village with a community of 501 people in which every two villagers exchange 100 personal coins with each other. As a result, each villager will have 50,000 coins of other villagers, while issuing and transferring 50,000 of their own coins to others, with the village achieving a total liquidity of 25,000,000 coins in circulation, solely based on mutual trust among villagers and without any external capital or credit. Such liquidity may jumpstart the village’s local economy, with the exposure of each villager to any other initially limited to 0.2% of the total credit they have issued.
Bank lending via bonds. The bank lends community coins to member $p$ in exchange for $p$-bonds with maturity and interest. For example, the bank mints 10,000 community coins for $p$, and $p$ mints 10,500 $p$-bonds maturing in six months—500 bonds as interest. The bank holds interest-bearing claims on $p$; $p$ holds liquid community coins accepted by all members. This is an instance of the loan instruments of Section 3 (items 2–4), with the bank as lender.
Deposits. A member may deposit community coins with the bank and receive community bonds maturing at $t$ with interest—for example, 10,000 community coins deposited, 10,200 community bonds maturing in six months returned. The bank’s income is the spread between lending and deposit rates.
Collateral. Personal coins—those acquired by community members through bilateral mutual credit—serve as collateral for bank loans. The bank may require borrower $p$ to deposit non-$p$-coins it holds into escrow (Section 4) as security. If $p$ defaults, the escrowed bonds are released to the bank. If $p$ fulfils the loan, the collateral is returned. The bank does not hold or manage personal coins directly; it holds $p$-bonds and, when needed, escrowed collateral.
Risk management. The bank manages risk through the maturity structure of its loan portfolio, collateral requirements at origination, the liquidity ratios of Section 5 applied to its borrowers, and the ability to sell debt at a discount (Section 3, item 6). As in traditional banking, a fixed-term loan cannot be recalled before maturity; the bank must arrange protection at origination or sell the debt to a third party.
Mutual currency recognition. An alternative to forming a community bank is for community members to mutually recognise each other’s currencies: $p$ and $q$ agree to accept each other’s coins as payment, each up to a per-issuer limit. This is a bilateral standing agreement that builds exposure organically as payments flow, without the upfront coin exchange of a credit line. If all community members mutually recognise each other’s coins, the community forms a currency area without a community bank. Similarly, if two banks mutually recognise each other’s currencies, their customers can transact across banks directly. This mechanism enables bottom-up formation of currency areas; its full treatment is deferred to the grassroots currencies paper (Shapiro, 2024b).
7. Implementation in GLP
Grassroots Logic Programs (GLP) (Shapiro, 2025) is a multiagent, concurrent, logic programming language designed for programming digital social contracts. A social contract is realised as a GLP program, which people load and activate voluntarily on their devices.
Here we implement the grassroots bonds transactions of Section 2—mint, swap, and redeem—as a GLP program, and demonstrate them through a simulated village economy involving six agents exercising the full gamut of financial instruments.
7.1. Bond Agent
The core of the implementation is the bond agent, a GLP process that manages each person’s holdings and executes transactions on their behalf. The agent maintains a list of bonds, a serial counter for minting, and communicates with other agents via typed message streams.
A bond is represented as bond(Issuer, Maturity, Serial). The core operations correspond directly to the formal specification of Section 2:
Mint. The agent creates $k$ bonds with a given maturity and adds them to its holdings:
  agent(Id, [msg(’_user’, Id1, mint(K, Maturity))|UserIn], ...) :-
      create_bonds(Id?, Maturity?, K?, NextSerial?, NewBonds),
      append(Holdings?, NewBonds?, NewHoldings),
      agent(Id?, UserIn?, ..., NewHoldings?, ...).
Swap. A swap is realised as a trade: the proposer specifies what they give and what they want as lists of lots lot(Issuer, Maturity, Count). The agent selects matching bonds from its holdings; if successful, it sends the bonds and an unbound response variable to the counterparty. The counterparty either accepts (binding the response to trade_accept(TheirBonds)) or declines (binding it to trade_decline(ReturnedBonds)). This two-phase protocol—propose then accept/decline—implements the volitional multiagent atomic transaction of Section 2.1: both parties must consent for the swap to commit.
The proposer side of the trade protocol:
  do_trade_result(ok, Id, Target, WantSpec, Selected,
      Remaining, UserIn, NetIn, Outs, NextSerial) :-
      lookup_send(friend(Target?),
          msg(Id?, Target?,
              trade_propose(WantSpec?, Selected?, TradeResp)),
          Outs?, Outs1),
      inject_trade_result(TradeResp?, Target?,
          UserIn?, UserIn1),
      agent(Id?, UserIn1?, NetIn?, Outs1?, Remaining?,
            NextSerial?).
The agent sends its selected bonds along with an unbound response variable TradeResp to the counterparty via the friend channel. Concurrently, inject_trade_result monitors TradeResp—when the counterparty binds it (to accept or decline), the result is injected into the agent’s input stream. This is the volitional multiagent atomic transaction: the unbound variable is the pending consent, and binding it commits or aborts the swap.
Note that the offered bonds are locked until the counterparty responds; if this is a risk that initiator of the swap does not wish to take, they can use a time-bound escrow agent instead.
Unification. All financial instruments—mutual credit, loans, pay, and redeem—are realised as combinations of mint and trade, with no dedicated commands or protocol machinery. Symmetric mutual credit is two mints followed by a trade; an asymmetric loan is a mint by the lender, a trade proposal, and a mint by the borrower who then accepts; pay is a trade with an empty want-specification. This unification simplifies the agent: the only operations are mint, trade, and escrow.
Trade classification. When the agent receives a trade proposal, it classifies it into one of three categories:
  classify_trade(_, [], _, payment).
  classify_trade(Id, _, [bond(I, 0, _)], redemption) :-
      Id? =?= I? | true.
  classify_trade(_, _, _, normal) :- otherwise | true.
Pay (no returned bonds) is auto-accepted—an agent cannot refuse accepting their own coins as payment. Redeem (exactly one of the agent’s own coins is offered) is also auto-accepted if the requested bond is available; otherwise the agent responds with a menu of available bonds—one per issuer, earliest maturity—so the proposer can retry an informed redemption request.
Coin redemption. In the formal model, when $p$ redeems a $q$-coin from $q$ it may specify any bond held by $q$, in effect assuming that $p$ knows what bonds $q$ holds. We approximate this in the GLP implementation as follows: if $q$ cannot respond with the requested bond, it reports to $p$ the type of bonds it holds (most mature of each type), allowing $p$ to choose from them. The caveat is that these bonds are not “locked” until $p$ decides what to do, allowing $q$, as irl, to “conceal assets”, e.g. by transferring them to another person or even to a Sybil identity controlled by $q$. While locking the reported assets until $p$ decides would avoid this problem, it also exposes $q$ to the risk of $p$ not responding, and even to coordinated attack, hence we chose not to do this. Still, high-risk/high-stakes transactions can use an escrow agent.
Escrow. The escrow mechanism is a concurrent race between two clauses:
  escrow(T, Bonds, _,
      escrow_bonds(Bonds?), escrow_expired) :-
      wait_until(T?) | true.
  escrow(_, Bonds, cancel,
      escrow_cancelled, escrow_bonds(Bonds?)).
The first clause suspends on a timer guard; the second suspends on the cancel signal. Whichever commits first determines the outcome—time-release or cancellation—with the other alternative abandoned. This employs GLP’s committed-choice nondeterminism: two concurrent alternatives race, and the first to satisfy its guard wins. The complete bond agent, including all helper procedures for bond selection, trade dispatch, redemption, and friend-channel management, is approximately 740 lines of typed GLP code, shown in Appendix B.
7.2. Village Market Demonstration
To demonstrate that the instruments of Sections 34 compose into a functioning economy, we implemented a simulated village market with six agents: Alice (baker), Bob (farmer), Charlie (carpenter), Diana (doctor), Eve (teacher), and Frank (fisherman). The scenario unfolds over a simulated month and exercises seven distinct financial operations:
  1. (1)
    Symmetric mutual credit. Alice and Bob each mint 15 coins and swap them 1-for-1, establishing mutual liquidity. Similarly for Alice–Charlie, Charlie–Eve, and Eve–Frank.
  2. (2)
    Asymmetric credit (loan). Diana lends Bob 20 diana-coins in exchange for 24 bob-bonds maturing on day 25—a 20% interest rate. Diana similarly lends Frank 15 coins for 18 bonds maturing day 28.
  3. (3)
    Pay. Bob pays Alice 5 alice-coins for bread; Eve pays Charlie 6 charlie-coins for a bookshelf; Frank pays Diana 3 diana-coins for a medical checkup.
  4. (4)
    Portfolio swap. Eve trades 5 frank-coins to Charlie for 5 alice-coins, acquiring the currency she needs to buy bread from Alice.
  5. (5)
    Escrow. Charlie hires Frank to build a dock and deposits 5 frank-coins in escrow for Frank with a 7-day cancellation window; Charlie does not cancel, the timer expires, and the coins are released to Frank as payment for building a dock.
  6. (6)
    Redeem. Frank redeems 5 diana-coins from Diana, reducing his exposure to Diana’s currency.
  7. (7)
    Sale of debt. Alice sells 10 bob-bonds to Eve for 4 frank-coins—a discounted sale reflecting Alice’s desire to reduce exposure to Bob’s debt.
All six agents run concurrently as GLP processes, communicating via typed message streams over a simulated 6-way network. Each agent emits a narrative trace describing its actions in natural language, alongside the raw GLP command/notification trace. Figure 1 shows the narrative output of all six agents.
Figure 1. Village Market: six-agent grassroots bond economy. Each panel shows one agent’s narrative trace. The scenario exercises symmetric mutual credit, asymmetric credit (loans with interest), pay, portfolio swap, escrow (time-release), redeem, and sale of debt. All agents run concurrently as GLP processes; the trace is produced by a running GLP program, the key part of it shown in Appendix B.
7.3. Implementation Summary
An intended smartphone deployment would include the bond agent (Appendix B,  740 lines of typed GLP code) and a UI; all other components are test infrastructure. The village market scenario of Section 7.2 is implemented by six scripted GLP actors totalling approximately 870 lines, each in its own module. The implementation is open-source and available at https://github.com/EShapiro2/GLP-Bonds.
The proof-of-concept represents each bond as a unit; a practical deployment would use range-denominated bonds to compress holdings of the same issuer and maturity into a single term.
8. Related Work
8.1. Global Cryptocurrencies
Global cryptocurrencies such as Bitcoin (Nakamoto and Bitcoin, 2008) and Ethereum (Buterin, 2014) are unbacked digital assets whose scarcity and transaction cost are controlled by the protocol. Grassroots coins (Shapiro, 2024b) and the grassroots bonds introduced here differ from global cryptocurrencies in their foundations, architecture, and the financial instruments they support. This section develops these contrasts along four axes: architectural design, source of value and liquidity, the nature of the financial instruments each system supports, and the attack surfaces each design exposes.
Architecture. Global cryptocurrencies maintain a single replicated ledger—a totally-ordered blockchain—via consensus among miners or validators. This architecture entails well-known costs: proof-of-work entails high-cost computing (Nakamoto and Bitcoin, 2008); proof-of-stake requires capital lockup (Wood, 2014); and all participants must obtain and maintain a copy of the global state. Transaction throughput is bounded by block size and block interval, and transaction fees arise from competition for inclusion in the next block.
Grassroots coins and bonds employ a fundamentally different architecture. Transaction cost is limited to operating a networked device. The system is grassroots (Shapiro, 2024a): multiple independent deployments can emerge without global resources and interoperate once interconnected.
The comparison table in (Lewis-Pye et al., 2023) details 18 dimensions along which the two architectures differ, including: one cryptocurrency per blockchain vs. one per person, mutually pegged; competitive block creation vs. decentralised and cooperative; global all-to-all dissemination vs. local grassroots dissemination; probabilistic finality vs. definite finality; and a business model based on ICOs and miner remuneration vs. a public good.
Source of value and liquidity. Global cryptocurrencies derive their value from speculative market demand; they are unbacked, and liquidity is obtained by mining or purchasing with external capital. Grassroots currencies are closer to ‘inside money’ (Cavalcanti and Wallace, 1999b)—a medium of exchange backed by private credit. Each person backs their coins and bonds with their offerings: goods and services for people and corporations; interest and transaction fees for banks; taxes for municipalities; and fiat currencies for central banks (Shapiro, 2024b). Liquidity arises not from capital injection but from mutual credit lines among persons that trust each other, and bond redemption—the obligation to redeem any coin one has issued against any bond one holds—pegs mutually-liquid currencies at a 1:1 exchange rate. Grassroots bonds extend this foundation by allowing credit to carry interest: lending liquid coins in exchange for future-maturity bonds, incentivising the provision of credit without requiring external capital.
Financial instruments on blockchains. The programmability of blockchains via smart contracts (Buterin, 2014; Sergey et al., 2019) has given rise to Decentralised Finance (DeFi) (Werner et al., 2022; Schär, 2021), which re-implements traditional financial instruments on-chain. Lending protocols (Gudgeon et al., 2020b; Leshner, 2020) offer overcollateralised loans with algorithmically-determined interest rates; decentralised exchanges provide token trading via automated market makers; and stablecoins offer price-stable media of exchange.
These DeFi instruments have analogues among the grassroots bond instruments of Sections 34: loans, credit lines, sale of debt, and options are all expressible as volitional swaps of grassroots bonds, without requiring a blockchain or consensus. Two differences are noteworthy. First, DeFi instruments require overcollateralisation enforced by smart contracts because counterparty identity is pseudonymous and trust is absent; grassroots instruments operate among persons with mutual trust, and collateral (Section 4) is available but not mandatory. Second, DeFi instruments incur gas fees for every state transition; grassroots instruments incur no fees beyond bilateral communication.
Crypto-native constructs. Three constructs are native to blockchain architecture and have no counterpart in grassroots bonds: flash loans, automated market makers (AMMs), and maximal extractable value (MEV). All three arise from properties of global consensus that grassroots systems lack by design. A comprehensive systematisation identifies 181 real-world DeFi attack incidents exploiting these properties (Zhou et al., 2023).
Flash loans (Qin et al., 2021) are uncollateralised loans that exist within a single atomic blockchain transaction: a borrower takes a loan, uses the funds across arbitrarily many smart contracts, and repays within the same transaction; if repayment fails, the entire transaction reverts. This is possible because blockchain transactions are atomic across the global state. Grassroots bonds have no global atomic transactions spanning multiple parties, and so flash loans cannot arise. This is a feature rather than a limitation: flash loans are the primary attack vector against DeFi protocols, with documented exploits yielding returns exceeding 500,000% (Qin et al., 2021). Their impossibility eliminates that entire class of vulnerability.
AMMs provide algorithmic price discovery via pooled liquidity and a bonding curve (e.g., the constant product formula $x\cdot y=k$). They require globally-shared mutable state—the liquidity pool—maintained by consensus. Grassroots bonds have no pooled global state; price discovery is via bilateral negotiation. The trade-off is that grassroots bonds lack automated price discovery but avoid impermanent loss, sandwich attacks (Daian et al., 2020), and the concentration of value in pool contracts that creates targets for exploitation.
MEV (Daian et al., 2020) is the profit that miners or validators can extract by reordering, inserting, or censoring transactions within a block. MEV manifests as frontrunning, backrunning, and sandwich attacks, and has been estimated to extract hundreds of millions of dollars annually from Ethereum users (Qin et al., 2022). MEV is an artefact of global transaction ordering. Grassroots bonds have no global ordering of transactions—each person’s blocklace fragment is independent—and so MEV does not exist.
Summary. Grassroots coins and bonds occupy a different point in the design space than global cryptocurrencies and their DeFi instruments. Where global cryptocurrencies achieve trustless operation through consensus at the cost of high transaction fees, capital requirements, and exposure to flash-loan attacks, MEV extraction, and sandwich attacks, grassroots bonds achieve consensus-free operation through bilateral trust at the cost of requiring social relationships for liquidity. The financial instruments are comparable in expressiveness (Sections 34), but differ in their trust assumptions, cost structure, and attack surface.
8.2. Finance
Here we survey additional related work: trust-based lending, informal community finance, digital social contracts, financial contract specification, credit networks, IOU-based currencies, mutual credit systems, tokenised bonds, off-chain payment protocols, and monetary theory.
Microfinance and peer-to-peer lending. The Grameen Bank model demonstrated that trust-based, collateral-free lending can work at scale via group liability (Morduch, 1999). Online peer-to-peer lending platforms extended this principle digitally: Lin et al. (Lin et al., 2013) showed that friendship networks on Prosper.com function as credible credit signals, increasing funding probability and predicting lower default rates; Iyer et al. (Iyer et al., 2016) found that peer lenders predict defaults 45% more accurately than credit scores. Grassroots bonds formalise this trust infrastructure into tradeable instruments: where microfinance relies on group liability and P2P platforms on centralised matching, grassroots bonds encode bilateral trust directly as bond swaps, with creditworthiness assessable via the liquidity measures of Section 5. This is relevant to financial inclusion: the World Bank Global Findex (Demirgüç-Kunt et al., 2022) estimates that 1.4 billion adults worldwide lack access to formal financial services, often in communities where mutual trust is abundant but capital is scarce—precisely the conditions under which grassroots bonds can bootstrap a local digital economy (Shapiro, 2024b).
Informal community finance. Rotating savings and credit associations (ROSCAs)—known as chit funds, tontines, or susus across cultures—are the dominant informal financial instrument in developing economies. Besley et al. (Besley et al., 1993) proved formally that both random and bidding ROSCAs yield higher utility than autarky by enabling earlier access to indivisible goods through community-based mutual commitment, without formal institutions. Ardener (Ardener, 1964) provided the foundational cross-cultural survey documenting their global prevalence. Hawala networks (El Qorchi et al., 2003) transfer value across borders on pure trust, without formal banking infrastructure. All three—ROSCAs, hawala, and grassroots bonds—share the same foundation: bilateral trust within a community substitutes for institutional intermediation. Grassroots bonds can be viewed as a digital formalisation that extends these informal mechanisms with the full expressiveness of a bond market.
Digital social contracts. Szabo (Szabo, 1997) introduced smart contracts as digitally-specified agreements with automated enforcement. Blockchain smart contracts realise this vision via consensus-bound execution. Cardelli et al. (Cardelli et al., 2020) introduced digital social contracts as an alternative: transition systems with agents performing digital speech acts, specified by a state-transducer language. Our definition refines theirs: a digital social contract is a voluntary agreement among persons, specified, fulfilled, and enforced digitally. Each financial instrument in Section 3 is a digital social contract—a volitional atomic transaction specifying conditions on the swap of grassroots bonds. Digital social contracts are thus the grassroots analogue of smart contracts: they achieve the same function (automated agreement enforcement) without blockchain infrastructure.
Financial contract specification. The foundational work on composable financial contract specification predates blockchain: Peyton Jones et al. (Peyton Jones et al., 2000) introduced a combinator library for describing financial contracts declaratively, inspiring subsequent blockchain domain-specific languages. Marlowe (Lamela Seijas and Thompson, 2018), Cardano’s financial contract DSL, specifies instruments declaratively with exhaustive static analysis; Scilla (Sergey et al., 2019) introduced automata-based contracts with formal safety guarantees. These languages demonstrate that financial instruments can be specified declaratively, yet they compile to consensus-bound smart contracts. Grassroots bonds take the same declarative approach—each instrument is a volitional atomic transaction (Section 2.1)—but enforce it bilaterally rather than via consensus, with each sovereign guaranteeing the security of transactions in their currency (Lewis-Pye et al., 2023).
Credit networks. Credit networks (Dandekar et al., 2011) study liquidity in networks of bilateral credit lines, with chain payments as the key instrument. Grassroots currencies inherit this structure: mutual credit lines formed via coin exchange, together with coin redemption, provide for chain payments as in credit networks (Shapiro, 2024b). Much of the body of knowledge regarding liquidity in credit networks carries over to grassroots bonds, with the caveat that lack of liquidity not only causes chain payment failures but also devalues the currencies of the liquidity-constrained parties involved.
IOU credit networks and personal currencies. Fugger’s proposal that money can be represented as IOUs in social trust networks (Fugger, 2004) is the earliest articulation of the idea underlying grassroots currencies. Ripple (Schwartz et al., 2014) implemented IOU-based credit routing on a blockchain, but re-introduced global consensus. SilentWhispers (Malavolta et al., 2017) proposed a distributed, privacy-preserving credit network that requires no ledger—the closest prior work to fully consensus-free financial transactions—but it addresses only payment routing, not the richer financial instruments that grassroots bonds support. Grassroots currencies (Shapiro, 2024b) implement Fugger’s vision without blockchain: personal coins are IOUs traded via bilateral transactions, with coin redemption pegging mutually-liquid currencies at 1:1. Grassroots bonds extend this by adding maturity dates, enabling interest-bearing credit.
Mutual credit and community currencies. Mutual credit systems are the closest precedent to grassroots bonds’ approach. Sardex, a mutual credit network in Sardinia, achieves significant real-economy impact; its cyclic transaction structure has been analysed in Nature Human Behaviour (Iosifidis et al., 2018). The Sarafu community currency in Kenya (Mattsson et al., 2022), operated on blockchain, is an academically-studied digital community currency deployment. Both systems create liquidity from bilateral trust without external capital—the same principle underlying grassroots bonds—but Sardex relies on a centralised operator and Sarafu on blockchain consensus.
Circles UBI (Circles, Retrieved 2021 https://joincircles.net/) implements blockchain-based personal currencies with a trust-network-mediated exchange rate, structurally similar to grassroots coin redemption among trusting parties; however, Circles enforces a uniform minting rate and relies on Ethereum miners for execution. The Trustlines Network (Hees,Heiko et al., 2021) extends mutual credit (LETS) principles to Ethereum, resulting in mutual credit lines similar to those of grassroots currencies but re-introducing blockchain consensus.
Tokenised bonds on blockchain. On-chain bond issuance represents the blockchain-native approach to the same problem grassroots bonds address. A Federal Reserve analysis (Carapella et al., 2023) examines tokenisation of real-world assets including the Santander (2019) and EIB (2021) bond issuances on Ethereum, identifying smart contract designs and financial stability implications. Security token offerings (STOs) extend this trend to debt instruments more broadly (Lambert et al., 2022). These tokenised bonds and STOs differ from grassroots bonds in three respects: they require a public blockchain and its associated gas costs; they rely on custodians to bridge on-chain tokens with off-chain legal obligations; and they are issued by institutions rather than by any person.
Traditional bond instruments. The loan structures of Section 3—zero-coupon, balloon, and fixed-payment loans—correspond to standard fixed-income instruments as classified by Fabozzi (Fabozzi and Fabozzi, 2021). The escrow-based instruments of Section 4—collateral, guarantees, and options—likewise have standard counterparts. Grassroots bonds differ from traditional bonds not in their financial structure but in three respects: they are issued by any person rather than by institutions, traded bilaterally rather than on exchanges, and enforced bilaterally rather than by legal contract. The economic effect is to democratise access to the bond market: any person can issue, hold, and trade bonds, with the community bank scenario of Section 6 illustrating how a village can achieve the financial functionality of a bank using only grassroots bonds.
Payment and state channels. Layer-2 scaling solutions reduce but do not eliminate reliance on blockchain consensus (Gudgeon et al., 2020a). Payment channels such as the Lightning Network (Poon and Dryja, 2016) enable off-chain bilateral transactions but fall back to the blockchain for dispute resolution. State channels (Dziembowski et al., 2018) generalise this to arbitrary state transitions. Boyen et al. (Boyen et al., 2016) proposed blockchain-free cryptocurrencies using a DAG structure, a prior attempt at the same design goal as grassroots currencies. Grassroots bonds’ bilateral transactions resemble state channel updates in structure, but differ fundamentally: state channels derive their security from the ability to publish the latest state on-chain in case of dispute, while grassroots bonds derive theirs from bilateral enforcement with no blockchain fallback. The atomic swap mechanism underlying grassroots bond transactions is related to cross-chain atomic swaps (Herlihy, 2018), which achieve trustless exchange across blockchains; grassroots bonds achieve the same atomicity bilaterally, without any blockchain.
Monetary theory. Grassroots bonds create inside money (Cavalcanti and Wallace, 1999b, a)—a medium of exchange backed by private credit rather than sovereign authority. Cavalcanti and Wallace (Cavalcanti and Wallace, 1999a) studied a population of bankers and nonbankers: bankers’ financial histories are public, enabling punishment for deviation, while nonbankers’ histories are private, making notes an essential medium of exchange. In grassroots currencies, everyone is between a banker and a nonbanker (Shapiro, 2024b): financial records are not public, but the holder of another person’s coins can choose which coin or bond to redeem against it among those held by the other person, requiring the coin holder to have access to such knowledge.
Kocherlakota’s result that “money is memory” (Kocherlakota, 1998) is directly relevant: grassroots bonds replace monetary tokens with cryptographic records of bilateral obligations, with agents holding the distributed memory that makes money possible. The Bank of England’s account of how commercial banks create money through lending (McLeay et al., 2014) describes a process structurally analogous to grassroots bond issuance: in both cases, the act of lending simultaneously creates both an asset (the loan) and a liability (the deposit or bond). The theoretical foundation for such endogenous money creation through credit is established by Palley (Palley, 2002). Brunnermeier et al. (Brunnermeier et al., 2019) provide a comprehensive framework connecting inside/outside money, token/account forms, and digital currency competition, within which grassroots bonds occupy a distinctive position as peer-issued, record-based inside money operating without either sovereign backing or blockchain consensus. Finally, Mundell’s theory of optimum currency areas (Mundell, 1961) is relevant to the regional bank and currency recognition mechanisms of Section 6, which allow grassroots currency areas to form and merge bottom-up rather than top-down.
9. Conclusion
Grassroots bonds extend grassroots coins (Shapiro, 2024b) with maturity dates, reframing coins—cash—as mature bonds. The entire specification reduces to two volitional transactions—mint and swap—from which the full gamut of financial instruments emerges as conditions on bond swaps, including credit lines, loans, sale of debt, forward contracts, options, and escrow-based instruments. Classical liquidity ratios apply to grassroots bonds, and the protocol is formally grassroots (Appendix A): independent deployments can emerge and interoperate without global resources. The formal specification was used by AI to derive a working implementation in GLP (Shapiro, 2025). A secure grassroots implementation of grassroots coins has been presented previously (Lewis-Pye et al., 2023). It applies directly to unilateral bond transactions (mint, pay, redeem); voluntary swaps between two parties employ bilateral atomic commitment, with escrow (Section 4) as a non-blocking alternative. Integrating these mechanisms in a secure implementation of GLP is a subject of ongoing work.
Appendix A Mathematical Foundations
This appendix collects the formal definitions underlying the grassroots framework. The definitions of configurations, transactions, active and stationary participants, and volitional transactions appear in Section 2.1; here we provide the additional machinery needed to define grassroots protocols and state the main theorem that the grassroots bonds protocol is grassroots.
A.1. Transition Systems
Definition A.1 (Transition System, Computation, Run).
A transition system is a tuple $TS=(S,s0,T)$, where $S$ is an arbitrary non-empty set of states, $s0\in S$ is the designated initial state, and $T\subset S^{2}$ is a set of transitions, where each transition $t\in T$ is a pair $(s,s^{\prime})$ of non-identical states $s\neq s^{\prime}\in S$, also written as $t=s\rightarrow s^{\prime}$. A computation of $TS$ is a (nonempty, potentially infinite) sequence of states $r=s_{1},s_{2},\cdots$ such that for every two consecutive states $s_{i},s_{i+1}\in r$, $s_{i}\rightarrow s_{i+1}\in T$. If $s_{1}=s0$ then the computation is called a run of $TS$.
Definition A.2 (Degree).
The degree of a transaction $t$ (unary, binary, …, $k$-ary) is the number of active participants in $t$, and the degree of a set of transactions $T$ is the maximal degree of any $t\in T$.
Definition A.3 (Multiagent Transition System).
Given agents $P\subset\Pi$ and an arbitrary set $S$ of local states with a designated initial local state $s0\in S$, a multiagent transition system over $P$ and $S$ is a transition system (Definition A.1) $TS=(C,c0,T)$ with configurations $C:=S^{P}$, initial configuration $c0:=\{s0\}^{P}$, and transitions $T\subseteq C^{2}$ being a set of transactions over $P$ and $S$.
A.2. Transaction Closure and Transactions-Based Systems
Definition A.4 (Transaction Closure).
Let $P\subset\Pi$, $S$ a set of local states, and $C:=S^{P}$. For a transaction $t=(c\rightarrow c^{\prime})$ over local states $S$ with participants $Q\subseteq P$, the $P$-closure of $t$, $t{\uparrow}P$, is the set of transitions over $P$ and $S$ defined by:
$t{\uparrow}P:=\{t^{\prime}\in C^{2}:\forall q\in Q.(t_{q}=t^{\prime}_{q})\wedge\forall p\in P\setminus Q.(p\text{ is stationary in }t^{\prime})\}$
If $R$ is a set of transactions, each $t\in R$ over some $Q\subseteq P$ and $S$, then the $P$-closure of $R$, $R{\uparrow}P:=\bigcup_{t\in R}t{\uparrow}P$.
Definition A.5 (Transactions-Based Multiagent Transition System).
Given agents $P\subset\Pi$, local states $S$ with initial local state $s0\in S$, and a set of transactions $R$, each $t\in R$ over some $Q\subseteq P$ and $S$, the transactions-based multiagent transition system over $P$, $S$, and $R$ is the multiagent transition system $TS=(S^{P},\{s0\}^{P},R{\uparrow}P)$.
A.3. Protocols and the Grassroots Property
A protocol is a family of multiagent transition systems, one for each set of agents $P\subset\Pi$, which share an underlying set of local states $\mathcal{S}$ with a designated initial state $s0$.
Definition A.6 (Local-States Function).
A local-states function $S:2^{\Pi}\to 2^{\mathcal{S}}$ maps every set of agents $P\subset\Pi$ to a set of local states $S(P)\subset\mathcal{S}$ that includes $s0$ and satisfies $P\subset P^{\prime}\subset\Pi\implies S(P)\subset S(P^{\prime})$.
Given a local-states function $S$, we use $C$ to denote configurations over $S$, with $C(P):=S(P)^{P}$ and $c0(P):=\{s0\}^{P}$.
Definition A.7 (Protocol).
A protocol $\mathcal{F}$ over a local-states function $S$ is a family of multiagent transition systems that has exactly one transition system $\mathcal{F}(P)=(C(P),c0(P),T(P))$ for every $P\subset\Pi$.
Definition A.8 (Projection).
Let $\emptyset\subset P\subset P^{\prime}\subset\Pi$. If $c^{\prime}$ is a configuration over $P^{\prime}$ then $c^{\prime}/P$, the projection of $c^{\prime}$ over $P$, is the configuration $c$ over $P$ defined by $c_{p}:=c^{\prime}_{p}$ for every $p\in P$.
Note that $c_{p}$, the state of $p$ in $c$, is in $S(P^{\prime})$, not in $S(P)$, and hence may include elements “alien” to $P$, e.g., bonds issued by $q\in P^{\prime}\setminus P$.
Definition A.9 (Oblivious, Interactive, Grassroots).
A protocol $\mathcal{F}$ is:
  1. (1)
    oblivious if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$, $T(P){\uparrow}P^{\prime}\subseteq T(P^{\prime})$.
  2. (2)
    interactive if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$ and every configuration $c\in C(P^{\prime})$ such that $c{/}P\in C(P)$, there is a computation $c\xrightarrow{*}c^{\prime}$ of $\mathcal{F}(P^{\prime})$ for which $c^{\prime}{/}P\notin C(P)$.
  3. (3)
    grassroots if it is oblivious and interactive.
Being oblivious implies that if a run of $\mathcal{F}(P^{\prime})$ reaches some configuration $c^{\prime}$, then anything $P$ could do on their own in the configuration $c^{\prime}/P$ (with a transition from $T(P)$), they can still do in the larger configuration $c^{\prime}$ (with a transition from $T(P^{\prime})$), effectively being oblivious to members of $P^{\prime}\setminus P$.
Being interactive is a weak liveness requirement: no matter what members of $P$ do, if they run within a larger set of agents it is always the case that they can eventually interact with non-$P$’s in a way that leaves “alien traces” in the local states of $P$, so that the resulting configuration $c^{\prime}/P$ could not have been produced by $P$ running on their own.
A.4. Transactions-Based Grassroots Protocols
Definition A.10 (Transactions Over a Local-States Function).
Let $S$ be a local-states function. A set of transactions $R$ is over $S$ if every transaction $t\in R$ is a multiagent transition over $Q$ and $S(Q^{\prime})$ for some $Q\subseteq Q^{\prime}\subset\Pi$. Given such a set $R$ and $P\subset\Pi$, $R(P):=\{t\in R:t\text{ is over }Q\text{ and }S(Q^{\prime}),Q\subseteq Q^{\prime}\subseteq P\}$.
Definition A.11 (Transactions-Based Protocol).
Let $S$ be a local-states function and $R$ a set of transactions over $S$. Then a protocol $\mathcal{F}$ over $R$ and $S$ includes for each set of agents $P\subset\Pi$ the transactions-based multiagent transition system $\mathcal{F}(P)$ over $P$, $S(P)$, and $R(P)$: $\mathcal{F}(P):=(S(P)^{P},\{s0\}^{P},R(P){\uparrow}P)$.
Definition A.12 (Interactive Transactions).
A set of transactions $R$ over a local-states function $S$ is interactive if for every $\emptyset\subset P\subset P^{\prime}\subset\Pi$ and every configuration $c\in C(P^{\prime})$ such that $c/P\in C(P)$, there is a computation $(c\xrightarrow{*}c^{\prime})\subseteq R(P^{\prime}){\uparrow}P^{\prime}$ for which $c^{\prime}/P\notin C(P)$.
Proposition A.13 ((Shapiro, 2026, Proposition 1)).
A transactions-based protocol is oblivious.
Theorem A.14 ((Shapiro, 2026, Theorem 1)).
A protocol over an interactive set of transactions is grassroots.
A.5. Grassroots Bonds Are Grassroots
Corollary A.15.
The grassroots bonds protocol (Definition 2.3) is grassroots.
The argument is straightforward: the grassroots bonds transactions are interactive. Given $P\subset P^{\prime}$ and a configuration $c$ in which all bonds held by members of $P$ are issued by members of $P$, a voluntary swap between $p\in P$ and $q\in P^{\prime}\setminus P$ produces $q$-bonds in $p$’s state—an alien trace that could not have been produced by $P$ running on their own.
We note that the GLP implementation (Section 7) extends the bonds transactions with the grassroots social graph (Shapiro, 2026) and direct messaging, as specified in (Shapiro, 2026) and implemented in (Shapiro, 2025). These extensions preserve the grassroots property: the social graph is itself grassroots (Shapiro, 2026), and direct messaging between friends does not alter the local states used for the grassroots argument.
Appendix B Bond Agent Source Code
The bond agent implementation consists of two GLP modules: self.glp, which defines all shared types and helper procedures visible to all modules via ancestor scoping; and agent.glp, which defines the bond agent and its local procedures. Together they comprise approximately 740 lines of typed GLP code. The full source, including the UI mediator, twelve test scenarios, and the six village market actors, is available at https://github.com/EShapiro2/GLP-Bonds.
Module 1: Shared Types and Helpers (self.glp)
%% self.glp — Bonds V2 shared type vocabulary
%% All protocol types for the Grassroots Bonds application.
%% Defined once here, visible to all modules via ancestor scoping.
-module(bonds).
-mode(system).
%% Stream(X) and Channel(In, Out) are predefined in prelude:
%%   Stream(X) ::= [] ; [X | Stream(X)].
%%   Channel(In, Out) ::= ch(In, Out?).
%% Bond representation: bond(Issuer, Maturity, Serial)
Bond ::= bond(Constant, Constant, Constant).
%% Trade types
Lot ::= lot(Constant, Constant, Constant).
TradeResponse ::= trade_accept(Stream(Bond))
               ; trade_decline(Stream(Bond))
               ; trade_decline_menu(Stream(Bond), Stream(Bond)).
%% Escrow cancel signal
EscrowCancel ::= cancel.
%% Escrow results
EscrowBenResult ::= escrow_bonds(Stream(Bond))
                  ; escrow_cancelled.
EscrowDepResult ::= escrow_bonds(Stream(Bond))
                  ; escrow_expired.
%% Friend messages (after connection established)
FriendContent ::= response(Response)
                ; text(Constant)
                ; trade_propose(Stream(Lot), Stream(Bond),
                    TradeResponse?)
                ; escrow_offer(Constant, EscrowBenResult).
FriendMsg     ::= msg(Constant, Constant, FriendContent).
FriendChannel ::= ch(Stream(FriendMsg),
                     Stream(FriendMsg)?).
Response ::= accept(FriendChannel) ; no.
Decision ::= yes ; no.
%% Network content (2-arg cold-call)
NetColdCall ::= intro(Constant, Response?).
%% User-to-agent message content
UserContent ::= mint(Constant, Constant)
  ; balance ; done
  ; connect(Constant)
  ; decision(Decision, Constant, PendingValue)
  ; response(Response)
  ; trade(Constant, Stream(Lot), Stream(Lot))
  ; accept_trade(Constant, PendingValue)
  ; reject_trade(Constant, PendingValue)
  ; deposit_escrow(Constant, Stream(Lot), Constant)
  ; cancel_escrow(Constant, PendingValue).
%% Pending values stored by mediator
PendingValue ::= response(Response?)
  ; trade_pending(TradeResponse?, Stream(Lot),
      Stream(Bond))
  ; escrow_pending(EscrowCancel?)
  ; error.
%% Agent-to-user message content
AgentContent ::= minted(Constant, Constant)
  ; balance_report(Stream(Bond))
  ; befriend(Constant, Response?)
  ; connected(Constant)
  ; rejected ; rejected(Constant)
  ; trade_proposed(Constant, Stream(Lot),
      TradeResponse?, Stream(Bond))
  ; trade_completed(Constant)
  ; trade_failed(Constant)
  ; trade_returned(Constant)
  ; escrow_deposited(Constant, Constant,
      EscrowCancel?)
  ; escrow_received(Constant, Constant)
  ; escrow_released(Constant)
  ; escrow_cancelled(Constant)
  ; escrow_expired(Constant)
  ; escrow_returned(Constant)
  ; escrow_failed(Constant)
  ; trade_returned_menu(Constant, Stream(Bond)).
AgentToUserMsg     ::= msg(Constant, Constant,
                           AgentContent).
MediatorToAgentMsg ::= msg(Constant, Constant,
                           UserContent).
%% Network input message
NetInMsg ::= msg(Constant, NetColdCall)
           ; msg(Constant, Constant, FriendContent).
%% Agent input stream (mediator + injected results)
UserInMsg ::= msg(Constant, Constant, UserContent)
  ; trade_complete(Constant, Stream(Bond))
  ; trade_returned_bonds(Constant, Stream(Bond))
  ; escrow_ben_released(Constant, Stream(Bond))
  ; escrow_ben_cancelled(Constant)
  ; escrow_dep_expired(Constant)
  ; escrow_dep_returned(Constant, Stream(Bond))
  ; trade_returned_bonds_menu(Constant,
      Stream(Bond), Stream(Bond)).
%% Output types (type union of agent + friend content)
OutputContent ::= AgentContent ; FriendContent.
OutputMsg ::= msg(Constant, Constant, OutputContent)
            ; msg(Constant, NetColdCall).
OutputKey   ::= ’_user’ ; ’_net’ ; friend(Constant).
OutputEntry ::= output(OutputKey, Stream(OutputMsg)?).
AgentChannel ::= ch(Stream(AgentToUserMsg),
                    Stream(MediatorToAgentMsg)?).
%% Shared helper procedures
procedure merge(Stream(X)?, Stream(X)?, Stream(X)).
merge([X|Xs], Ys, [X?|Zs?]) :-
    merge(Ys?, Xs?, Zs).
merge(Xs, [Y|Ys], [Y?|Zs?]) :-
    merge(Xs?, Ys?, Zs).
merge([], Ys, Ys?).
merge(Xs, [], Xs?).
procedure lookup_send(OutputKey?, OutputMsg?,
    Stream(OutputEntry)?, Stream(OutputEntry)).
lookup_send(Key, Msg, Outs, Outs1?) :-
    ground(Key?) |
    lookup_send_step(Key?, Msg?, Outs?, Outs1).
procedure lookup_send_step(OutputKey?, OutputMsg?,
    Stream(OutputEntry)?, Stream(OutputEntry)).
lookup_send_step(Key, Msg,
    [output(K, [Msg?|Out1?])|Rest],
    [output(K?, Out1)|Rest?]) :-
    Key? =?= K? | true.
lookup_send_step(Key, Msg,
    [output(K, Out?)|Rest],
    [output(K?, Out)|Rest1?]) :-
    otherwise |
    lookup_send_step(Key?, Msg?, Rest?, Rest1).
lookup_send_step(_, _, [], []).
procedure add_output(OutputKey?, Stream(OutputMsg),
    Stream(OutputEntry)?, Stream(OutputEntry)).
add_output(Key, Out?, Outs, [output(Key?, Out)|Outs?]).
procedure close_outputs(Stream(OutputEntry)?).
close_outputs([output(_, [])|Outs]) :-
    close_outputs(Outs?).
close_outputs([]).
procedure inject_msg(Response?, Constant?, Constant?,
    Stream(UserInMsg)?, Stream(UserInMsg)).
inject_msg(Resp, Target, Id, Ys,
    [msg(Target?, Id?, response(Resp?))|Ys?]) :-
    known(Resp?) | true.
inject_msg(Resp, Target, Id, [Y|Ys], [Y?|Ys1?]) :-
    inject_msg(Resp?, Target?, Id?, Ys?, Ys1).
procedure create_bonds(Constant?, Constant?,
    Constant?, Constant?, Stream(Bond)).
create_bonds(Issuer, Maturity, K, Serial,
    [bond(Issuer?, Maturity?, Serial?)|Rest?]) :-
    K? > 0 |
    K1 := K? - 1, S1 := Serial? + 1,
    create_bonds(Issuer?, Maturity?, K1?, S1?, Rest).
create_bonds(_, _, 0, _, []).
procedure append(Stream(Bond)?, Stream(Bond)?,
    Stream(Bond)).
append([X|Xs], Ys, [X?|Zs?]) :-
    append(Xs?, Ys?, Zs).
append([], Ys, Ys?).
procedure select_bonds_exact(Constant?, Constant?,
    Constant?, Stream(Bond)?, Constant,
    Stream(Bond), Stream(Bond)).
select_bonds_exact(_, _, 0, Hs, ok, [], Hs?).
select_bonds_exact(Issuer, Maturity, K,
    [bond(I, M, S)|Rest], Status?,
    [bond(I?, M?, S?)|Sel?], Rem?) :-
    K? > 0, Issuer? =?= I?, Maturity? =?= M? |
    K1 := K? - 1,
    select_bonds_exact(Issuer?, Maturity?, K1?,
        Rest?, Status, Sel, Rem).
select_bonds_exact(Issuer, Maturity, K, [B|Rest],
    Status?, Sel?, [B?|Rem?]) :-
    otherwise |
    select_bonds_exact(Issuer?, Maturity?, K?,
        Rest?, Status, Sel, Rem).
select_bonds_exact(_, _, K, [], fail, [], []) :-
    K? > 0 | true.
procedure select_bonds_by_spec(Stream(Lot)?,
    Stream(Bond)?, Constant,
    Stream(Bond), Stream(Bond)).
select_bonds_by_spec(
    [lot(Issuer, Maturity, Count)|Rest], Holdings,
    Status?, AllSel?, FinalRem?) :-
    ground(Issuer?), ground(Maturity?),
    ground(Count?) |
    select_bonds_exact(Issuer?, Maturity?, Count?,
        Holdings?, LotStatus, LotSel, LotRem),
    select_by_spec_continue(LotStatus?, Rest?,
        LotSel?, LotRem?,
        Status, AllSel, FinalRem).
select_bonds_by_spec([], Holdings, ok, [],
    Holdings?).
procedure select_by_spec_continue(Constant?,
    Stream(Lot)?, Stream(Bond)?, Stream(Bond)?,
    Constant, Stream(Bond), Stream(Bond)).
select_by_spec_continue(ok, Rest, LotSel, LotRem,
    Status?, AllSel?, FinalRem?) :-
    select_bonds_by_spec(Rest?, LotRem?,
        Status, RestSel, FinalRem),
    append(LotSel?, RestSel?, AllSel).
select_by_spec_continue(fail, _, LotSel, LotRem,
    fail, LotSel?, LotRem?).
procedure bind_trade_accept(TradeResponse,
    Stream(Bond)?).
bind_trade_accept(trade_accept(Bonds?), Bonds).
procedure bind_trade_decline(TradeResponse,
    Stream(Bond)?).
bind_trade_decline(trade_decline(Bonds?), Bonds).
procedure inject_trade_result(TradeResponse?,
    Constant?, Stream(UserInMsg)?,
    Stream(UserInMsg)).
inject_trade_result(trade_accept(TheirBonds), From,
    Ys, [trade_complete(From?, TheirBonds?)|Ys?]) :-
    ground(From?), ground(TheirBonds?) | true.
inject_trade_result(trade_decline(OurBonds), From,
    Ys,
    [trade_returned_bonds(From?, OurBonds?)|Ys?]) :-
    ground(From?), ground(OurBonds?) | true.
inject_trade_result(
    trade_decline_menu(OurBonds, Menu), From, Ys,
    [trade_returned_bonds_menu(From?, OurBonds?,
        Menu?)|Ys?]) :-
    ground(From?), ground(OurBonds?),
    ground(Menu?) | true.
inject_trade_result(Resp, From, [Y|Ys],
    [Y?|Ys1?]) :-
    inject_trade_result(Resp?, From?, Ys?, Ys1).
procedure classify_trade(Constant?, Stream(Lot)?,
    Stream(Bond)?, Constant).
classify_trade(_, [], _, payment).
classify_trade(Id, _, [bond(I, 0, _)], redemption) :-
    Id? =?= I? | true.
classify_trade(_, _, _, normal) :-
    otherwise | true.
procedure build_menu(Constant?, Stream(Bond)?,
    Stream(Bond)).
build_menu(Id, Holdings, Menu?) :-
    build_menu_acc(Id?, Holdings?, [], Menu).
procedure build_menu_acc(Constant?, Stream(Bond)?,
    Stream(Bond)?, Stream(Bond)).
build_menu_acc(Id, [bond(I, _, _)|Rest], Acc,
    Menu?) :-
    Id? =?= I? |
    build_menu_acc(Id?, Rest?, Acc?, Menu).
build_menu_acc(Id, [bond(I, M, S)|Rest], Acc,
    Menu?) :-
    otherwise |
    menu_update(I?, M?, S?, Acc?, Acc1),
    build_menu_acc(Id?, Rest?, Acc1?, Menu).
build_menu_acc(_, [], Acc, Acc?).
procedure menu_update(Constant?, Constant?,
    Constant?, Stream(Bond)?, Stream(Bond)).
menu_update(I, M, S, [bond(I2, M2, _)|Rest],
    [bond(I2?, M?, S?)|Rest?]) :-
    I? =?= I2?, M? < M2? | true.
menu_update(I, _, _, [bond(I2, M2, S2)|Rest],
    [bond(I2?, M2?, S2?)|Rest?]) :-
    I? =?= I2? | true.
menu_update(I, M, S, [B|Rest], [B?|Rest1?]) :-
    otherwise |
    menu_update(I?, M?, S?, Rest?, Rest1).
menu_update(I, M, S, [], [bond(I?, M?, S?)]).
procedure escrow(Constant?, Stream(Bond)?,
    EscrowCancel?, EscrowBenResult, EscrowDepResult).
escrow(T, Bonds, _,
    escrow_bonds(Bonds?), escrow_expired) :-
    wait_until(T?) | true.
escrow(_, Bonds, cancel,
    escrow_cancelled, escrow_bonds(Bonds?)).
procedure inject_escrow_ben_result(EscrowBenResult?,
    Constant?, Stream(UserInMsg)?,
    Stream(UserInMsg)).
inject_escrow_ben_result(escrow_bonds(Bonds), From,
    Ys,
    [escrow_ben_released(From?, Bonds?)|Ys?]) :-
    ground(From?), ground(Bonds?) | true.
inject_escrow_ben_result(escrow_cancelled, From, Ys,
    [escrow_ben_cancelled(From?)|Ys?]) :-
    ground(From?) | true.
inject_escrow_ben_result(Resp, From, [Y|Ys],
    [Y?|Ys1?]) :-
    inject_escrow_ben_result(Resp?, From?,
        Ys?, Ys1).
procedure inject_escrow_dep_result(EscrowDepResult?,
    Constant?, Stream(UserInMsg)?,
    Stream(UserInMsg)).
inject_escrow_dep_result(escrow_bonds(Bonds), Target,
    Ys,
    [escrow_dep_returned(Target?, Bonds?)|Ys?]) :-
    ground(Target?), ground(Bonds?) | true.
inject_escrow_dep_result(escrow_expired, Target, Ys,
    [escrow_dep_expired(Target?)|Ys?]) :-
    ground(Target?) | true.
inject_escrow_dep_result(Resp, Target, [Y|Ys],
    [Y?|Ys1?]) :-
    inject_escrow_dep_result(Resp?, Target?,
        Ys?, Ys1).
procedure bind_escrow_cancel(EscrowCancel).
bind_escrow_cancel(cancel).
procedure lookup_pending(ReqId?, PendingValue,
    Stream(PendingEntry)?, Stream(PendingEntry)).
lookup_pending(ReqId, Val?,
    [pending(Id, Val) | Rest], Rest?) :-
    ReqId? =?= Id? | true.
lookup_pending(ReqId, Val?, [P | Rest],
    [P? | Rest1?]) :-
    otherwise |
    lookup_pending(ReqId?, Val, Rest?, Rest1).
lookup_pending(_, error, [], []).
Module 2: Bond Agent (agent.glp)
%% agent.glp — Bond agent module (v2)
%% Exports: agent/6
%% Types and shared helpers: from self.glp
-module(agent).
-mode(system).
%% Merge — local copy (Stream(X) subtyping limitation)
procedure merge(Stream(X)?, Stream(X)?, Stream(X)).
merge([X|Xs], Ys, [X?|Zs?]) :-
    merge(Ys?, Xs?, Zs).
merge(Xs, [Y|Ys], [Y?|Zs?]) :-
    merge(Xs?, Ys?, Zs).
merge([], Ys, Ys?).
merge(Xs, [], Xs?).
%% Response handling — local (calls merge)
procedure bind_response(Decision?, Constant?,
    Response, Stream(OutputEntry)?,
    Stream(OutputEntry), Stream(NetInMsg)?,
    Stream(NetInMsg)).
bind_response(yes, From, accept(RetCh?),
    Outs, Outs1?, In, In1?) :-
    new_channel(RetCh, LocalCh) |
    handle_response(accept(LocalCh?), From?,
        Outs?, Outs1, In?, In1).
bind_response(no, _, no, Outs, Outs1?, In, In?) :-
    lookup_send(’_user’,
        msg(agent, ’_user’, rejected), Outs?, Outs1).
procedure handle_response(Response?, Constant?,
    Stream(OutputEntry)?, Stream(OutputEntry),
    Stream(NetInMsg)?, Stream(NetInMsg)).
handle_response(accept(ch(FIn, FOut?)), From,
    Outs, Outs2?, In, In1?) :-
    ground(From?) |
    add_output(friend(From?), FOut, Outs?, Outs1),
    lookup_send(’_user’,
        msg(agent, ’_user’, connected(From?)),
        Outs1?, Outs2),
    merge(In?, FIn?, In1).
handle_response(no, From, Outs, Outs1?, In, In?) :-
    ground(From?) |
    lookup_send(’_user’,
        msg(agent, ’_user’, rejected(From?)),
        Outs?, Outs1).
%% Trade — select bonds then send or fail
procedure do_trade(Constant?, Constant?,
    Stream(Lot)?, Stream(Lot)?, Stream(Bond)?,
    Stream(UserInMsg)?, Stream(NetInMsg)?,
    Stream(OutputEntry)?, Constant?).
do_trade(Id, Target, GiveSpec, WantSpec, Holdings,
    UserIn, NetIn, Outs, NextSerial) :-
    select_bonds_by_spec(GiveSpec?, Holdings?,
        Status, Selected, Remaining),
    do_trade_result(Status?, Id?, Target?,
        WantSpec?, Selected?, Remaining?,
        UserIn?, NetIn?, Outs?, NextSerial?).
procedure do_trade_result(Constant?, Constant?,
    Constant?, Stream(Lot)?, Stream(Bond)?,
    Stream(Bond)?, Stream(UserInMsg)?,
    Stream(NetInMsg)?, Stream(OutputEntry)?,
    Constant?).
do_trade_result(ok, Id, Target, WantSpec, Selected,
    Remaining, UserIn, NetIn, Outs, NextSerial) :-
    lookup_send(friend(Target?),
        msg(Id?, Target?,
            trade_propose(WantSpec?, Selected?,
                TradeResp)),
        Outs?, Outs1),
    inject_trade_result(TradeResp?, Target?,
        UserIn?, UserIn1),
    agent(Id?, UserIn1?, NetIn?, Outs1?,
        Remaining?, NextSerial?).
do_trade_result(fail, Id, Target, _, Selected,
    Remaining, UserIn, NetIn, Outs, NextSerial) :-
    append(Selected?, Remaining?, OrigHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’, trade_failed(Target?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        OrigHoldings?, NextSerial?).
%% Fill trade or decline
procedure handle_trade_fill(Constant?, Constant?,
    Constant?, TradeResponse, Stream(Bond)?,
    Stream(Bond)?, Stream(Bond)?,
    Stream(UserInMsg)?, Stream(NetInMsg)?,
    Stream(OutputEntry)?, Constant?).
handle_trade_fill(ok, Id, From,
    trade_accept(Selected?), OfferedBonds, Selected,
    Remaining, UserIn, NetIn, Outs, NextSerial) :-
    append(Remaining?, OfferedBonds?, NewHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’, trade_completed(From?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NextSerial?).
handle_trade_fill(fail, Id, From,
    trade_decline(OfferedBonds?), OfferedBonds,
    Selected, Remaining, UserIn, NetIn, Outs,
    NextSerial) :-
    append(Selected?, Remaining?, OrigHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’, trade_failed(From?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        OrigHoldings?, NextSerial?).
%% Trade dispatch — auto-accept vs. user decision
procedure trade_dispatch(Constant?, Constant?,
    Constant?, Stream(Lot)?, Stream(Bond)?,
    TradeResponse, Stream(Bond)?,
    Stream(UserInMsg)?, Stream(NetInMsg)?,
    Stream(OutputEntry)?, Constant?).
trade_dispatch(payment, Id, From, _, OfferedBonds,
    trade_accept([]), Holdings, UserIn, NetIn,
    Outs, NextSerial) :-
    append(Holdings?, OfferedBonds?, NewHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’, trade_completed(From?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NextSerial?).
trade_dispatch(redemption, Id, From, WantSpec,
    OfferedBonds, TradeResp?, Holdings, UserIn,
    NetIn, Outs, NextSerial) :-
    select_bonds_by_spec(WantSpec?, Holdings?,
        Status, Selected, Remaining),
    redemption_result(Status?, Id?, From?,
        TradeResp, OfferedBonds?, Selected?,
        Remaining?, UserIn?, NetIn?, Outs?,
        NextSerial?).
trade_dispatch(normal, Id, From, WantSpec,
    OfferedBonds, TradeResp?, Holdings, UserIn,
    NetIn, Outs, NextSerial) :-
    lookup_send(’_user’,
        msg(agent, ’_user’,
            trade_proposed(From?, WantSpec?,
                TradeResp, OfferedBonds?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        Holdings?, NextSerial?).
%% Redemption result
procedure redemption_result(Constant?, Constant?,
    Constant?, TradeResponse, Stream(Bond)?,
    Stream(Bond)?, Stream(Bond)?,
    Stream(UserInMsg)?, Stream(NetInMsg)?,
    Stream(OutputEntry)?, Constant?).
redemption_result(ok, Id, From,
    trade_accept(Selected?), OfferedBonds, Selected,
    Remaining, UserIn, NetIn, Outs, NextSerial) :-
    append(Remaining?, OfferedBonds?, NewHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’, trade_completed(From?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NextSerial?).
redemption_result(fail, Id, From,
    trade_decline_menu(OfferedBonds?, Menu?),
    OfferedBonds, Selected, Remaining,
    UserIn, NetIn, Outs, NextSerial) :-
    append(Selected?, Remaining?, OrigHoldings),
    redemption_reject(Id?, From?, OrigHoldings?,
        Menu, UserIn?, NetIn?, Outs?, NextSerial?).
procedure redemption_reject(Constant?, Constant?,
    Stream(Bond)?, Stream(Bond),
    Stream(UserInMsg)?, Stream(NetInMsg)?,
    Stream(OutputEntry)?, Constant?).
redemption_reject(Id, From, OrigHoldings, Menu?,
    UserIn, NetIn, Outs, NextSerial) :-
    ground(OrigHoldings?) |
    build_menu(Id?, OrigHoldings?, Menu),
    lookup_send(’_user’,
        msg(agent, ’_user’, trade_failed(From?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        OrigHoldings?, NextSerial?).
%% Escrow — select bonds and create escrow or fail
procedure do_deposit_escrow(Constant?, Constant?,
    Stream(Lot)?, Constant?, Stream(Bond)?,
    Stream(UserInMsg)?, Stream(NetInMsg)?,
    Stream(OutputEntry)?, Constant?).
do_deposit_escrow(Id, Target, GiveSpec, ReleaseTime,
    Holdings, UserIn, NetIn, Outs, NextSerial) :-
    select_bonds_by_spec(GiveSpec?, Holdings?,
        Status, Selected, Remaining),
    do_deposit_escrow_result(Status?, Id?, Target?,
        ReleaseTime?, Selected?, Remaining?,
        UserIn?, NetIn?, Outs?, NextSerial?).
procedure do_deposit_escrow_result(Constant?,
    Constant?, Constant?, Constant?, Stream(Bond)?,
    Stream(Bond)?, Stream(UserInMsg)?,
    Stream(NetInMsg)?, Stream(OutputEntry)?,
    Constant?).
do_deposit_escrow_result(ok, Id, Target, ReleaseTime,
    Selected, Remaining, UserIn, NetIn, Outs,
    NextSerial) :-
    ground(Selected?) |
    escrow(ReleaseTime?, Selected?, CancelSignal?,
        BenResult, DepResult),
    inject_escrow_dep_result(DepResult?, Target?,
        UserIn?, UserIn1),
    lookup_send(friend(Target?),
        msg(Id?, Target?,
            escrow_offer(ReleaseTime?, BenResult?)),
        Outs?, Outs1),
    lookup_send(’_user’,
        msg(agent, ’_user’,
            escrow_deposited(Target?, ReleaseTime?,
                CancelSignal)),
        Outs1?, Outs2),
    agent(Id?, UserIn1?, NetIn?, Outs2?,
        Remaining?, NextSerial?).
do_deposit_escrow_result(fail, Id, Target, _,
    Selected, Remaining, UserIn, NetIn, Outs,
    NextSerial) :-
    append(Selected?, Remaining?, OrigHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’, escrow_failed(Target?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        OrigHoldings?, NextSerial?).
%% Bond agent — exported entry point
exported procedure agent(Constant?,
    Stream(UserInMsg)?, Stream(NetInMsg)?,
    Stream(OutputEntry)?, Stream(Bond)?, Constant?).
%% Mint
agent(Id,
    [msg(’_user’, Id1, mint(K, Maturity))|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    Id? =?= Id1?, ground(K?), ground(Maturity?) |
    NewNextSerial := NextSerial? + K?,
    create_bonds(Id?, Maturity?, K?, NextSerial?,
        NewBonds),
    append(Holdings?, NewBonds?, NewHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’, minted(K?, Maturity?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NewNextSerial?).
%% Balance
agent(Id, [msg(’_user’, Id1, balance)|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    Id? =?= Id1?, ground(Holdings?) |
    lookup_send(’_user’,
        msg(agent, ’_user’,
            balance_report(Holdings?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        Holdings?, NextSerial?).
%% Done
agent(Id, [msg(’_user’, Id1, done)|_], _, Outs,
    _, _) :-
    Id? =?= Id1? | close_outputs(Outs?).
%% Connect (cold call)
agent(Id,
    [msg(’_user’, Id1, connect(Target))|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    Id? =?= Id1?, ground(Target?) |
    lookup_send(’_net’,
        msg(Target?, intro(Id?, Resp)),
        Outs?, Outs1),
    inject_msg(Resp?, Target?, Id?,
        UserIn?, UserIn1),
    agent(Id?, UserIn1?, NetIn?, Outs1?,
        Holdings?, NextSerial?).
%% Decision on cold-call
agent(Id, [msg(’_user’, Id1,
    decision(Dec, From, response(Resp?)))|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    Id? =?= Id1? |
    bind_response(Dec?, From?, Resp,
        Outs?, Outs1, NetIn?, NetIn1),
    agent(Id?, UserIn?, NetIn1?, Outs1?,
        Holdings?, NextSerial?).
%% Cold-call response (injected)
agent(Id,
    [msg(From, Id1, response(Resp))|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    Id? =?= Id1? |
    handle_response(Resp?, From?,
        Outs?, Outs1, NetIn?, NetIn1),
    agent(Id?, UserIn?, NetIn1?, Outs1?,
        Holdings?, NextSerial?).
%% Cold-call from network
agent(Id, UserIn,
    [msg(Id1, intro(From, Resp?))|NetIn],
    Outs, Holdings, NextSerial) :-
    Id? =?= Id1? |
    lookup_send(’_user’,
        msg(agent, ’_user’, befriend(From?, Resp)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        Holdings?, NextSerial?).
%% Trade command
agent(Id, [msg(’_user’, Id1,
    trade(Target, GiveSpec, WantSpec))|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    Id? =?= Id1?, ground(Target?),
    ground(GiveSpec?), ground(WantSpec?) |
    do_trade(Id?, Target?, GiveSpec?, WantSpec?,
        Holdings?, UserIn?, NetIn?, Outs?,
        NextSerial?).
%% Trade complete (injected — accept)
agent(Id,
    [trade_complete(From, TheirBonds)|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    ground(Id?), ground(From?),
    ground(TheirBonds?) |
    append(Holdings?, TheirBonds?, NewHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’, trade_completed(From?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NextSerial?).
%% Trade returned (injected — decline)
agent(Id,
    [trade_returned_bonds(From, OurBonds)|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    ground(Id?), ground(From?),
    ground(OurBonds?) |
    append(Holdings?, OurBonds?, NewHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’, trade_returned(From?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NextSerial?).
%% Trade returned with menu
agent(Id, [trade_returned_bonds_menu(From,
    OurBonds, Menu)|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    ground(Id?), ground(From?),
    ground(OurBonds?), ground(Menu?) |
    append(Holdings?, OurBonds?, NewHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’,
            trade_returned_menu(From?, Menu?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NextSerial?).
%% Incoming trade_propose
agent(Id, UserIn, [msg(From, Id1,
    trade_propose(WantSpec, OfferedBonds,
        TradeResp?))|NetIn],
    Outs, Holdings, NextSerial) :-
    Id? =?= Id1?, ground(From?),
    ground(WantSpec?), ground(OfferedBonds?) |
    classify_trade(Id?, WantSpec?,
        OfferedBonds?, TradeClass),
    trade_dispatch(TradeClass?, Id?, From?,
        WantSpec?, OfferedBonds?, TradeResp,
        Holdings?, UserIn?, NetIn?, Outs?,
        NextSerial?).
%% Accept trade
agent(Id, [msg(’_user’, Id1,
    accept_trade(From,
        trade_pending(TradeResp?, WantSpec,
            OfferedBonds)))|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    Id? =?= Id1?, ground(From?),
    ground(WantSpec?), ground(OfferedBonds?) |
    select_bonds_by_spec(WantSpec?, Holdings?,
        Status, Selected, Remaining),
    handle_trade_fill(Status?, Id?, From?,
        TradeResp, OfferedBonds?, Selected?,
        Remaining?, UserIn?, NetIn?, Outs?,
        NextSerial?).
%% Reject trade
agent(Id, [msg(’_user’, Id1,
    reject_trade(From,
        trade_pending(TradeResp?, _,
            OfferedBonds)))|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    Id? =?= Id1?, ground(From?),
    ground(OfferedBonds?) |
    bind_trade_decline(TradeResp, OfferedBonds?),
    agent(Id?, UserIn?, NetIn?, Outs?,
        Holdings?, NextSerial?).
%% Deposit escrow
agent(Id, [msg(’_user’, Id1,
    deposit_escrow(Target, GiveSpec,
        ReleaseTime))|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    Id? =?= Id1?, ground(Target?),
    ground(GiveSpec?), ground(ReleaseTime?) |
    do_deposit_escrow(Id?, Target?, GiveSpec?,
        ReleaseTime?, Holdings?, UserIn?, NetIn?,
        Outs?, NextSerial?).
%% Escrow expired (injected)
agent(Id, [escrow_dep_expired(Target)|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    ground(Id?) |
    lookup_send(’_user’,
        msg(agent, ’_user’,
            escrow_expired(Target?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        Holdings?, NextSerial?).
%% Escrow returned (injected)
agent(Id,
    [escrow_dep_returned(Target, OurBonds)|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    ground(Id?), ground(Target?),
    ground(OurBonds?) |
    append(Holdings?, OurBonds?, NewHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’,
            escrow_returned(Target?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NextSerial?).
%% Cancel escrow
agent(Id, [msg(’_user’, Id1,
    cancel_escrow(Target,
        escrow_pending(CancelSignal?)))|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    Id? =?= Id1?, ground(Target?) |
    bind_escrow_cancel(CancelSignal),
    agent(Id?, UserIn?, NetIn?, Outs?,
        Holdings?, NextSerial?).
%% Incoming escrow offer
agent(Id, UserIn, [msg(From, Id1,
    escrow_offer(ReleaseTime, BenResult))|NetIn],
    Outs, Holdings, NextSerial) :-
    Id? =?= Id1?, ground(From?),
    ground(ReleaseTime?) |
    inject_escrow_ben_result(BenResult?, From?,
        UserIn?, UserIn1),
    lookup_send(’_user’,
        msg(agent, ’_user’,
            escrow_received(From?, ReleaseTime?)),
        Outs?, Outs1),
    agent(Id?, UserIn1?, NetIn?, Outs1?,
        Holdings?, NextSerial?).
%% Escrow released (injected)
agent(Id,
    [escrow_ben_released(From, Bonds)|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    ground(Id?), ground(From?), ground(Bonds?) |
    append(Holdings?, Bonds?, NewHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’,
            escrow_released(From?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NextSerial?).
%% Escrow cancelled (injected)
agent(Id, [escrow_ben_cancelled(From)|UserIn],
    NetIn, Outs, Holdings, NextSerial) :-
    ground(Id?) |
    lookup_send(’_user’,
        msg(agent, ’_user’,
            escrow_cancelled(From?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        Holdings?, NextSerial?).
%% Termination
agent(_, [], _, Outs, _, _) :-
    close_outputs(Outs?).